A wind turbine is rotating counterclockwise at 0.5 rev/s and slows to a stop in 10 s. Its blades are 20 m in length


1)The angular acceleration of the turbine is


2)The centripetal acceleration of the tip of the blades at t=0s is


3)The magnitude of the total linear acceleration of the tip of the blades at t=0s is


4)The direction of the total linear acceleration of the tip of the blades at t=0s is Blank ____° counterclockwise


1
Expert's answer
2021-03-18T16:52:52-0400

1) We can find the angular acceleration of the turbine as follows:


"\omega=\omega_0+\alpha t,""\alpha=\dfrac{\omega-\omega_0}{t},""\alpha=\dfrac{0-0.5\ \dfrac{rev}{s}\cdot\dfrac{2\pi\ rad}{1\ rev}}{10\ s}=-0.314\ \dfrac{rad}{s^2}."

The sign minus means that the turbine decelerates.

2) We can find centripetal acceleration of the tip of the blades as follows:


"a_c=\omega_0^2R=(0.5\ \dfrac{rev}{s}\cdot\dfrac{2\pi\ rad}{1\ rev})^2\cdot20\ m=197.2\ \dfrac{m}{s^2}."

3) Let's first find the tangential acceleration of the tip of the blades:


"a_t=\alpha R=(-0.314\ \dfrac{rad}{s^2})\cdot20\ m=-6.28\ \dfrac{m}{s^2}."

Finally, we can find the magnitude of the total linear acceleration of the tip of the blades from the Pythagorean theorem:


"a=\sqrt{a_c^2+a_t^2}=\sqrt{(197.2\ \dfrac{m}{s^2})^2+(-6.28\ \dfrac{m}{s^2})^2}=197.3\ \dfrac{m}{s^2}."

4) We can find the direction of the total linear acceleration of the tip of the blades from the geometry:


"tan\theta=\dfrac{|a_t|}{|a_c|},""\theta=tan^{-1}(\dfrac{|a_t|}{|a_c|}),""\theta=tan^{-1}(\dfrac{|-6.28\ \dfrac{m}{s^2}|}{|197.3\ \dfrac{m}{s^2}|})=1.82^{\circ}."